Consider the decomposition of



6x3+x−3/x2−2x+1



into partial fractions

1
Expert's answer
2022-06-08T17:06:44-0400

Use long division to find the quotient and remainder of "\frac{6x^3+x-3}{x^2-2x+1}" and then rewrite it as the quotient plus the remainder over the denominator:




Now that we have that "\frac{6x^3+x-3}{x^2-2x+1}=(6x+12)+\frac{19x+15}{x^2-2x+1}". We have only to find the partial fraction decomposition of "\frac{19x-15}{x^2-2x+1}".


Express as a ratio of expanded polynomials in lowest terms: "\frac{19x-15}{x^2-2x+1}=\frac{19x-15}{(x-1)^2}".


The partial fraction expansion is of the form: "\frac{19x-15}{(x-1)^2}=\frac{a}{x-1}+\frac{b}{(x-1)^2}".


Multiply both sides by "(x-1)^2" and simplify: "19x-15=a(x-1)+b".


Expand and collect in terms of powers of "x": "19x-15=-a+bx+c".


Equate coefficients on both sides, yielding 2 equation in 2 unknows:

"\begin{cases}
-15=b-a\\19=a
\end{cases}"


The solution: "a=19", "b=4".


Therefore: "\frac{19x-15}{x^2-2x+1}=\frac{4}{(x-1)^2}+\frac{19}{x-1}".


Compiling our results, we have that "\frac{6x^3+x-3}{x^2-2x+1}=12+6x+\frac{4}{(x-1)^2}+\frac{19}{x-1}".